OpenStudy (anonymous):

if y^2=sin(x), determine dy/dx

jimthompson5910 (jim_thompson5910):

y^2=sin(x) derive both sides with respect to x to get d/dx[y^2]=d/dx[sin(x)] 2y*y' = cos(x) Then solve for y' to get 2y*y' = cos(x) y' = cos(x)/2y y' = cos(x)/(2*sqrt(sin(x))) Finally, replace y' with dy/dx dy/dx = cos(x)/(2*sqrt(sin(x)))

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np